OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (1, 5) and is parallel to the line y = –3x – 3?

OpenStudy (anonymous):

y = -3x + 8

OpenStudy (anonymous):

my possible answer are: thank you for helping me 3x + y = 6 3x + y = 8 x + 3y = –6 x – 3y = 0

OpenStudy (anonymous):

so it will be: 3x+y=8

OpenStudy (anonymous):

3x + y = 8 is the same thing as y = -3x + 8

OpenStudy (anonymous):

thank you..i just got it